# Confidential Prompting: Protecting User Prompts from Cloud LLM Providers

### v2

@@ 4.1 Prefill in CVM
Upon receiving Alice’s $n$-token prompt, Ted computes the attention states $K_{\mathrm{Pvt}}\in\mathbb{R}^{n\times d}$ and $V_{\mathrm{Pvt}}\in\mathbb{R}^{n\times d}$, and sends the first token to Eve. The LLM weights are discarded after prefill. We introduce two optimizations. (i) Since LLM weights are used only once, we employ a memory-efficient strategy where Ted fetches the weights as needed and discards them after use, limiting memory usage to a single layer. (ii) Not all portions of the user prompt are sensitive; for instance, common system messages can be public. Users can designate sensitive parts of the prompt using the <confidential/> tag. Ted stores $K_{\mathrm{Pvt}}$ and $V_{\mathrm{Pvt}}$ only for the sensitive parts, migrating the rest to $K_{\mathrm{Pub}}$ and $V_{\mathrm{Pub}}$ on Eve.
Security analysis of model confidentiality  Ted requires read-only access to Eve’s LLM weights in the prefill phase. As assumed in §3.2, Eve can inspect Ted’s outbound data to prevent weight leakage to Alice. Eve is able to check the integrity of the outbound data because Eve knows the LLM responses and the responses are not encrypted. We note that even the use of PO (§5) does not require Ted to send Alice the value of the index of the authentic response among the $\lambda+1$ candidates, which would create a side channel for the leakage of LLM weights. This is because Alice shares the key (or called seed) of the PRF with Ted(§3.2) and, as a result, already knows the index.
4.2 Secure Multi-party Decoding (SMD)
@@
We formulate the decoding operation as a secure two-party computation based on the online softmax calculation [39], allowing Eve to retrieve the full attention output without knowing $K_{\mathrm{Pvt}}$ and $V_{\mathrm{Pvt}}$. Specifically, we introduce the following theorem for two-party decoding:
@@
Theorem 4.1 is lossless and its proof is available in the appendix. This theorem serves as the foundation for the design of SMD, which offers two key benefits. First, Eve can batch process all public states ($Q,K_{\mathrm{Pub}},V_{\mathrm{Pub}}$) of different users in parallel using a single set of LLM weights. Second, computations in Ted have a fixed cost and do not involve LLM weights, allowing Ted to retain only a small amount of memory for the private states. In addition, SMD only induces a constant communication overhead of $12d+$ 1 floating points in each round. We note that computing $\gamma_{\mathrm{Pvt}}$ and $\gamma_{\mathrm{Pub}}$ individually is numerically unstable due to its exponential term. Instead, we compute $\gamma_{\mathrm{Pvt}}=\sum\text{exp}(QK_{\mathrm{Pvt}}^{\top}-m_{\mathrm{Pvt}})$, where $m_{\mathrm{Pvt}}=\max(QK_{\mathrm{Pvt}}^{\top})$. The coefficients in Theorem 4.1 become $\frac{\gamma_{\mathrm{Pvt}}}{\gamma_{\mathrm{Pvt}}+\alpha\gamma_{\mathrm{Pub}}}$ and $\frac{\gamma_{\mathrm{Pub}}}{\alpha^{-1}\gamma_{\mathrm{Pvt}}+\gamma_{\mathrm{Pub}}}$, where $\alpha=\exp(m_{\mathrm{Pub}}-m_{\mathrm{Pvt}})$.
@@ 4.3 Security Analysis
Under the assumptions outlined in §3.2, Alice’s prompt remains secure both during transmission to Ted and throughout the prefill phase within the CVM. During the decoding phase, Ted ensures that neither the prompt nor the private KV cache is exposed to Eve. In other words, maintaining the private KV cache within the CVM ensures prompt confidentiality. Even if Eve disregards the private KV cache stored in the CVM and attempts to recompute the KV attention states, she is unable to do so because the prompt itself remains secure within the CVM as well. Instead, Eve only has access to the generated tokens and the output of $\sigma(Q_{\mathrm{New}}K_{\mathrm{Pvt}}^{\top})V_{\mathrm{Pvt}}$ from Ted, which can not be directly reversed to retrieve the prompt or private KV attention states.
This security analysis suggests that Eve has the only opportunity to extract the private user prompt (or the embedding of the prompt $X_{\mathrm{Pvt}}\in\mathbb{R}^{n_{\mathrm{Pvt}}\times d}$, where $K_{\mathrm{Pvt}}=X_{\mathrm{Pvt}}W_{K}$ and $V_{\mathrm{Pvt}}=X_{\mathrm{Pvt}}W_{V}$) through (semi) black-box reconstruction attacks. In a simpler case, Eve can attempt to find an approximate solution $\tilde{X}_{\mathrm{Pvt}}$ for
$\lVert\mathrm{Ted}(Q)-\sigma(Q(\tilde{X}_{\mathrm{Pvt}}W_{K})^{T})(\tilde{X}_{\mathrm{Pvt}}W_{V})\rVert=0$ (2)
where $\mathrm{Ted}(Q)$ is the response from Ted. Eve can also perform a stronger reconstruction attack using public KV attention states of the generated tokens:
$\lVert\mathrm{Eve}(Q)-\sigma(Q\times\text{concat}(\tilde{X}_{\mathrm{Pvt}}W_{K},K_{\mathrm{Pub}})^{T})$
@@
where $\mathrm{Eve}(Q)$ is the result from Theorem 4.1. Reconstruction attacks generally require a large number of observations to be effective. However, because of the autoregressive nature of LLMs, they can theoretically generate outputs of unlimited length, resulting in an unlimited number of public states. Although relying solely on SMD could secure the user prompt against typical adversaries who might compromise the cloud, it leaves the system vulnerable to adversaries with unlimited memory and computational resources.
@@ 5 Prompt Obfuscation
We present prompt obfuscation (PO), a novel defense against reconstruction attacks. Inspired by chaffing and winnowing [30, 31], PO aims to fundamentally limit Eve’s ability to accurately observe Alice’s prompts, at the cost of generating redundant tokens. As mentioned in §2.3.1, the premise of confidentiality without encryption in Rivest’s work is that (1) an eavesdropper cannot distinguish between authentic and fake messages based on their content, while (2) the receiver can. PO achieves (1) by leveraging LLMs’ ability to perform infilling and (2) by exploiting the protocol’s reflective nature, in which Ted sends intermediate data to Eve for processing and the output is reflected back to Ted. During the prefill phase in the CVM (§4.1), Ted creates a set of virtual prompts by replacing segments of the authentic user prompt with sampled fake n-grams. In the decoding phase (§4.2), Ted instructs Eve to decode all virtual prompts and the authentic prompt in parallel, while keeping track of the authentic prompt’s index to relay the authentic output back to Alice. Figure 4 provides an example of PO.
@@
1:User’s prompt $S$ as token sequence, target subsequence $S_{\text{target}}$, parameters $\epsilon$ and $\lambda_{\text{max}}$. LM is a pre-trained language model, and $\{F_{i}\}$ are the associated pre-defined control sequences to format the prompt. The + operator indicates concatenation, while $S_{a}-S_{b}$ indicates removing subsequence $S_{b}$ from $S_{a}$.
@@
4:$\tilde{S}\leftarrow Format(\{F_{i}\},S_{\text{context}},n)$triangleright Format $S$ for the infilling task.
@@
11:    for each $C\in S_{fake}$ do $\triangleright$ Find all tokens with probability within the same quantized bin as $\rho$.
12:         $X\sim\text{LM}(\tilde{S}+C)$triangleright Each “bin” has a width of $\epsilon^{\prime}$. There are $\lceil\frac{1}{\epsilon^{\prime}}\rceil$ bins in total.
13:         $S_{\text{new}}\leftarrow S_{\text{new}}\cup\{C+x\mid\left\lfloor\frac{\rho}{\epsilon^{\prime}}\right\rfloor\cdot\epsilon^{\prime}<P(X=x)\leq\left(\left\lfloor\frac{\rho}{\epsilon^{\prime}}\right\rfloor+1\right)\cdot\epsilon^{\prime}\}$
@@
15:    $S_{fake}\leftarrow\text{top-}\lambda_{\text{max}}(S_{\text{new}})$triangleright Only select $\lambda_{\text{max}}$ candidates to keep the compute cost static.
@@ 5.1 Sampling Virtual Prompts
We introduce Greedy Quantized Sampling (GQS), a simple algorithm designed to sample fake n-grams based on the user’s prompt $S$, as illustrated in Algorithm 1. We ask users to mark sensitive subsequences in the prompts with <redacted/> tags. For simplicity, we assume that the sensitive subsequences within a prompt are independent and process them separately, denoted $S_{target}$. The target subsequence $S_{target}$ may represent a name, location, medical condition, or other sensitive information. The goal of GQS is to generate multiple fake replacement subsequences that appear as authentic as the original $S_{target}$. In addition to $S$ and $S_{target}$, GQS requires parameters $\epsilon$ and $\lambda_{max}$, as well as a pre-trained language model $LM$. For brevity, we denote the algorithm as $GQS(S,S_{target},\epsilon)$.
Algorithm 1 leverages the ability of LLMs to infill missing subsequence in a prompt, that is, replacing $S_{target}$ with $S$. Some LLMs are specifically trained for infilling tasks [40, 41, 42]. It has also been shown that simple autoregressive LLMs can perform infilling when the prompts are appropriately formatted [43], such as using the $Format$ function to obtain the formatted prompt $\tilde{S}$ in Algorithm 1. Sequences $\{F_{i}\}$ refer to predefined control sequences used to format the prompt for the infilling task, which are determined by the chosen language model $LM$. For example, tokens like <prefix> and <suffix> are commonly used in fill-in-the-middle (FIM) models [44].
There are two key parameters in GQS that control the number of fake n-grams to generate: $0<\epsilon\leq 1$ and $\lambda_{\text{max}}\in\mathbb{N}$. $\epsilon$ is the error bound for the fake n-grams, while $\lambda_{\text{max}}$ controls the maximum number of candidates to consider at each step. These parameters can be adjusted according to security requirements and Ted’s available memory. Given the prompt $S$ and the target subsequence $S_{\text{target}}$ from Alice, Ted samples fake n-grams $S_{\text{fake}}$ using $GQS(S,S_{target},\epsilon)$ and constructs a set of virtual prompts by replacing $S_{\text{target}}$ with each $C\in S_{\text{fake}}$. Alice may also provide an additional parameter, $\lambda_{min}$. If fewer than $\lambda_{min}$ virtual prompts are generated, Ted will stop in the prefill phase and alert Alice to a potential information leak if the process continues.
@@
5.2 Compact Representation of Virtual Prompts
We introduce a memory-efficient approach to store virtual prompts within CVMs. This approach significantly reduces memory overhead while introducing a small approximation error to the decoding of virtual prompts. It does not affect the final accuracy, as the decoding of the authentic prompt remains unaffected, and the virtual ones are eventually discarded.
Given $\lambda$ virtual prompts, instead of storing all $K_{\mathrm{Pvt}}$ and $V_{\mathrm{Pvt}}$ for each virtual prompt, which would require memory proportional to $\lambda|S|$, we compute the attention states for the shared part $S-S_{\text{target}}$ separately, and share them across all virtual prompts, but not for the authentic prompt. This allows Ted to store attention states only for the individual fake n-grams $C\in S_{\text{fake}}$, reducing the memory cost to store all virtual prompts from $\lambda|S|$ to $|S-S_{\text{target}}|+\lambda|S_{\text{target}}|$.
This method introduces a small approximation error for the decoding of virtual prompts due to the attention masking effect, since the attention states for the shared $S-S_{\text{target}}$ are computed without the fake replacement $C\in S_{\text{fake}}$. However, this optimization does not affect the final accuracy, as the authentic prompt is correctly decoded.
@@ 5.3 Security Analysis
PO adds an extra layer of protection to SMD (§4), defending against (semi) black-box reconstruction attacks in two folds. (1) First, it linearly increases the cost of reconstruction attacks, while incurring only sublinear resource usage within the CVM (§5.2) and minimal performance degradation (§7.2). (2) Second, even in the worst case where an adversary reconstructs all $\lambda+1$ prompts (with $\lambda$ being the number of virtual prompts), the adversary’s probability of correctly guessing the authentic prompt $S$ improves by only $\epsilon$ beyond randomly guessing from the $\lambda+1$ possible prompts, assuming that the language model $LM$ accurately captures the true distribution $\mathcal{D}$. We next analyze (2) in more detail.
@@
Given a user prompt $S$ and a subsequence $S_{target}$, let $S_{fake}=\text{GQS}(S,S_{\text{target}},\epsilon)$ and $\Delta=\Delta_{U}(\mathcal{D},\mathcal{D}_{\text{LM}})$, where $U=\{S_{target}\}\cup S_{fake}$. Then $\{S_{target}\}$ and $S_{fake}$ are $(\epsilon+2\Delta)$-close over $\mathcal{D}$.
Theorem 5.1 implies that if both $\epsilon$ and $\Delta$ are small, the virtual prompts generated as described in §5.1 would appear as authentic as the user prompt $S$. More precisely, the maximum difference between the probabilities of any pair of authentic and fake subsequences is $\epsilon+2\Delta$, given $S$ as context. The proof is provided in the appendix. Suppose that we have an authentic prompt $S$ and a set of virtual prompts $S_{1},S_{2},\ldots,S_{\lambda}$. Let $\mathcal{D}_{\text{LM}}$ be the distribution of all possible sequences modeled by $LM$. Theorem 5.1 essentially states that an adversary with knowledge of $\mathcal{D}$ can guess the authentic prompt $S$ with a probability $\epsilon+2\Delta$ better than randomly choosing from $\lambda+1$ possible prompts. Therefore, the security of PO can be characterized by three factors: (1) the set size, $\lambda$; (2) the error bound of sampling, $\epsilon$; and (3) the gap between the true distribution and its approximation, $\Delta=\Delta_{U}(\mathcal{D},\mathcal{D}_{\text{LM}})$.
@@
However, PO does not fully guarantee confidentiality when the true distribution $\mathcal{D}$ differs from $\mathcal{D}_{\text{LM}}$. This can happen when $\mathcal{D}_{\text{LM}}$ is simply not sufficiently sophisticated. In such cases, PO serves as an obfuscation, as its name suggests, increases the cost of cryptanalysis, but does not fully prevent it. Furthermore, the above analysis suggests that an adversary with greater knowledge of the true distribution $\mathcal{D}$ gains an advantage in guessing. However, it is important to note that $\mathcal{D}$ is not always accessible to adversaries, meaning that the actual error bound in practice is likely smaller than our theoretical analysis.
Another scenario in which PO does not provide full confidentiality is when the prompt $S$ is unique. In such cases, it may be impossible to generate fake n-grams without using a large $\epsilon$ or small $\lambda$, regardless of whether $\mathcal{D}$ is well captured by $\mathcal{D}_{\text{LM}}$ or not. To prevent a potential information leak, PO allows users to configure the desired $\epsilon$ and $\lambda_{min}$. If $\lambda<\lambda_{min}$, Ted will stop in the prefill phase and alert Alice.

### v3

@@ 4.1 Prefill in CVM
Upon receiving the $n$-token prompt from user, the CVM computes the attention states $K_{\mathrm{Pvt}}\in\mathbb{R}^{n\times d}$ and $V_{\mathrm{Pvt}}\in\mathbb{R}^{n\times d}$, and sends the first token to the cloud LLM. The LLM weights are discarded after prefill. We introduce an optimization to further reduce the CVM memory footprint: Since LLM weights are used only once, we employ a memory-efficient strategy where the CVM fetches the weights on demand and discards them after use, limiting memory usage to a single layer.
Security analysis of model confidentiality  The CVM requires read-only access to the cloud LLM weights in the prefill phase. As mentioned in §3.2, the cloud vendor can inspect outbound data from the CVM, preventing LLM weights leakage to the user. The cloud LLM is able to check the integrity of the outbound data because it is the cloud LLM which generates the output tokens, knowing the exact value of the responses.
We note that even the use of PO (§5) does not require the CVM to send any encrypted data to the user during prefill and decode phases, otherwise forming a potential side channel to leak LLM weights. In particular, both the user and its CVM know the index of the authentic response among the $\lambda+1$ candidates, without any communication after CVM initialization. This is because they share the key (or called seed) of the PRF and thus both already know the index in advance.
4.2 Secure Partitioned Decoding (SPD)
@@
We formulate the decoding as a secure two-party computation based on the online softmax calculation [36], allowing the cloud LLM to retrieve the full attention score $Y$ without knowing $K_{\mathrm{Pvt}}$ and $V_{\mathrm{Pvt}}$. Specifically, we introduce the following theorem for two-party decoding:
@@
Theorem 4.1 is lossless and its proof is available in the appendix. This theorem serves as the foundation for the SPD design, which offers two key benefits. First, the cloud LLM can batch process all public states ($Q,K_{\mathrm{Pub}},V_{\mathrm{Pub}}$) of different users in parallel using a single set of LLM weights. Second, computations in CVMs have a fixed cost and do not involve LLM weights, allowing the CVMs to retain only a small amount of memory for the private states. In addition, SPD only induces a constant communication overhead of $12d+$ 1 floating points in each round. We note that computing $\gamma_{\mathrm{Pvt}}$ and $\gamma_{\mathrm{Pub}}$ individually is numerically unstable due to its exponential term. Instead, we compute $\gamma_{\mathrm{Pvt}}=\sum\text{exp}(QK_{\mathrm{Pvt}}^{\top}-m_{\mathrm{Pvt}})$, where $m_{\mathrm{Pvt}}=\max(QK_{\mathrm{Pvt}}^{\top})$. The coefficients in Theorem 4.1 become $\frac{\gamma_{\mathrm{Pvt}}}{\gamma_{\mathrm{Pvt}}+\alpha\gamma_{\mathrm{Pub}}}$ and $\frac{\gamma_{\mathrm{Pub}}}{\alpha^{-1}\gamma_{\mathrm{Pvt}}+\gamma_{\mathrm{Pub}}}$, where $\alpha=\exp(m_{\mathrm{Pub}}-m_{\mathrm{Pvt}})$.
@@ 4.3 Security Analysis
User prompts and the private KV cache are confidential throughout their lifetime. Under the assumptions outlined in §3.2, user prompts are secured during transmission to the CVM. They remain in the CVM without any exposure in prefill phase. During decode, the CVM exposes only the private attention score $A_{pvt}$, ensuring neither the prompt nor the private KV cache is directly exposed. The cloud LLM only has access to the generated tokens and the private attention score $A_{pvt}$ received from the CVM, which can not be directly reversed to uncover the prompt or the private KV cache.
This security analysis implies an attacker can uncover the private user prompt (or the embedding of the prompt $X_{\mathrm{Pvt}}\in\mathbb{R}^{n_{\mathrm{Pvt}}\times d}$, where $K_{\mathrm{Pvt}}=X_{\mathrm{Pvt}}W_{K}$ and $V_{\mathrm{Pvt}}=X_{\mathrm{Pvt}}W_{V}$) only via (semi) black-box reconstruction attacks. In a simpler case, the attacker can attempt to find an approximation of $\tilde{X}_{\mathrm{Pvt}}$ for
$\lVert A_{pvt}-\sigma(Q(\tilde{X}_{\mathrm{Pvt}}W_{K})^{T})(\tilde{X}_{\mathrm{Pvt}}W_{V})\rVert=0.$ (2)
An attacker can also perform a stronger reconstruction attack using public KV cache of the generated tokens:
$\lVert Y-\sigma(Q\times\text{concat}(\tilde{X}_{\mathrm{Pvt}}W_{K},K_{\mathrm{Pub}})^{T})$
@@
where $Y$ is the result from Theorem 4.1. It is worth noting that, an attempt to uncover a prompt directly from its known response is actually a reconstruction attack, which is usually not as effective as subsection 4.3, because the latter contains more information, i.e., the distribution of the next token, than just the final output tokens.
Reconstruction attacks generally require a large number of observations to be effective. However, due to the autoregressive nature of LLMs, they can theoretically generate unlimited output tokens, resulting in a large number of public states.
To summarize, relying solely on SPD could secure the user prompt against typical adversaries who compromise the cloud but have limited computational resources. But it leaves the system vulnerable to adversaries who have unlimited computational resources to perform reconstruction attacks.
@@ 5 Prompt Obfuscation
We present Prompt Obfuscation (PO), a novel defense against reconstruction attacks. Inspired by chaffing and winnowing [48, 5], PO aims to limit cloud LLM’s ability to accurately observe user prompts, at the cost of generating redundant tokens. The premise of chaffing and winnowing is that (1) an eavesdropper cannot distinguish between authentic and fake messages based on their content, while (2) the receiver can. PO achieves (1) by leveraging LLMs’ ability to perform infilling and (2) by exploiting the protocol’s reflective nature, in which the CVM sends intermediate data to the cloud LLM for processing and the output is reflected back to the CVM.
To obfuscate attackers and protect sensitive information in a user prompt, the CVM first identifies and tags sensitive segments in the prompt. Users can deploy any existing tools like Google’s PII detection or train a small language model (SLM) for tagging automatically. Then the CVM generates $\lambda$ virtual prompts by replacing the sensitive segments with sampled fake n-grams. Figure 14 in §A.3 is an example. These $\lambda$ virtual prompts and their private KV cache always reside in the CVM, in the same way as the authentic one. The CVM and the cloud LLM collaborate to decode all $\lambda+1$ prompts in parallel, with the cloud LLM being unable to distinguish the virtual prompts from the authentic one. Only the user and its CVM know which prompt and response are authentic.
@@
1:User’s prompt $S$ as token sequence, target subsequence $S_{\text{target}}$, parameters $\epsilon$ and $\lambda_{\text{max}}$. LM is a pre-trained language model, and $\{Ctr_{i}\}$ are the associated pre-defined control sequences to format the prompt. The + operator indicates concatenation, while $S_{a}-S_{b}$ indicates removing subsequence $S_{b}$ from $S_{a}$.
@@
4:$\tilde{S}\leftarrow Format(\{Ctr_{i}\},S_{\text{context}},n)$triangleright Format $\tilde{S}$ for the infilling task.
@@
11:    for each $S_{fake}\in F$ do $\triangleright$ Find all tokens with probability within the same quantized bin as $\rho$.
12:         $X\sim\text{LM}(\tilde{S}+S_{fake})$triangleright Each “bin” has a width of $\epsilon^{\prime}$. There are $\lceil\frac{1}{\epsilon^{\prime}}\rceil$ bins in total.
13:         $F_{\text{new}}\leftarrow F_{\text{new}}\cup\{S_{fake}+x\mid\left\lfloor\frac{\rho}{\epsilon^{\prime}}\right\rfloor\cdot\epsilon^{\prime}<P(X=x)\leq\left(\left\lfloor\frac{\rho}{\epsilon^{\prime}}\right\rfloor+1\right)\cdot\epsilon^{\prime}\}$
@@
15:    $F\leftarrow\text{top-}\lambda_{\text{max}}(F_{\text{new}})$triangleright Only select $\lambda_{\text{max}}$ candidates to keep the compute cost static.
@@ 5.1 Sampling Virtual Prompts
We introduce Greedy Quantized Sampling (GQS), a simple algorithm designed to sample fake n-grams based on the user’s prompt $S$, as illustrated in Algorithm 1. We assume the automatic tagging tool marks sensitive segments, i.e., sensitive token subsequences, with <redacted/> tags. For simplicity, we assume the sensitive subsequences within a prompt are independent and the CVM processes them one by one, denoted $S_{target}$. The target subsequence $S_{target}$ may represent a name, location, medical condition, or other sensitive information. The goal of GQS is to generate multiple fake replacement subsequences that appear as authentic as the original $S_{target}$. In addition to $S$ and $S_{target}$, GQS requires parameters $\epsilon$ and $\lambda_{max}$, as well as a pre-trained language model $LM$. For brevity, we denote the algorithm as $GQS(S,S_{target},\epsilon)$.
Algorithm 1 leverages the ability of LLMs to infill missing subsequence in a prompt, that is, replacing $S_{target}$ with $S_{fake}\in F$. Some LLMs are trained for infilling tasks [11, 12, 4]. It has also been shown that simple autoregressive LLMs can perform infilling when the prompts are appropriately formatted [39], such as using the $Format$ function to obtain the formatted prompt $\tilde{S}$ in Algorithm 1. Sequences $\{Ctr_{i}\}$ refer to predefined control sequences used to format the prompt, which are determined by the chosen language model $LM$. For example, tokens like <prefix> and <suffix> are commonly used in fill-in-the-middle (FIM) models [28].
There are two key parameters in GQS that control the number of fake n-grams to generate: $0<\epsilon\leq 1$ and $\lambda_{\text{max}}\in\mathbb{N}$. $\epsilon$ is the error bound for the fake n-grams, while $\lambda_{\text{max}}$ controls the maximum number of candidates to consider at each step. These parameters adjustable according to security requirements and the CVM’s available memory. Given the prompt $S$ and the target subsequence $S_{\text{target}}$, the CVM samples a set of fake n-grams $F$ using $GQS(S,S_{target},\epsilon)$ and constructs a set of virtual prompts by replacing $S_{\text{target}}$ with $S_{\text{fake}}\in F$. The user may provide an additional parameter, $\lambda_{min}$. If fewer than $\lambda_{min}$ virtual prompts are generated, the CVM will stop and alert the user to a potential information leak if it continues.
@@
Compact Representation of Virtual Prompts  We introduce a memory-efficient approach to store virtual prompts within CVMs. This approach significantly reduces memory overhead while introducing a small approximation error to the decoding of virtual prompts. It does not affect the final accuracy, as the decoding of the authentic prompt remains unaffected, and the virtual ones are eventually discarded.
Given $\lambda$ virtual prompts, instead of storing all $K_{\mathrm{Pvt}}$ and $V_{\mathrm{Pvt}}$ for each virtual prompt, which would require memory proportional to $\lambda|S|$, we compute the attention states for the shared part $S-S_{\text{target}}$ separately, and share them across all virtual prompts, but not for the authentic prompt. This allows the CVM to store attention states only for the individual fake n-grams $S_{\text{fake}}\in F$, reducing the memory cost to store all virtual prompts from $\lambda|S|$ to $|S-S_{\text{target}}|+\lambda|S_{\text{target}}|$.
@@ 5.2 Security Analysis
In §3.3, we analyze SPD and PO are complementary and collaboratively defend against attacks that would likely succeed if either technique were employed independently. In this section, we analyze in the worst case where an attacker reconstructs all $\lambda+1$ prompts, the probability of correctly guessing the authentic prompt is $\epsilon+1/(\lambda+1)$, assuming the language model $LM$ accurately captures the true distribution $\mathcal{D}$.
Security Model  The key of analyzing PO is how to quantify “authenticity” of virtual prompts. To achieve this, we introduce the concept of $\epsilon$-close over a distribution $\mathcal{D}$ to bound the “distance” between two sets of messages:
Definition 5.1 ($\epsilon$-close over $\mathcal{D}$).
$\mathcal{D}$ refers to the true language distribution, which may not be known to the user or any party. The language model $LM$ used by the user to create the fake n-grams can approximate $\mathcal{D}$, denoted as $\mathcal{D_{\text{LM}}}$. We define the “distance” between two distributions over a set of messages $U$ as
If $\Delta_{U}(\mathcal{D},\mathcal{D_{\text{LM}}})$ is close to zero, we say $LM$ provides a sufficiently accurate approximation of $\mathcal{D}$ over subdomain $U$.
Intuitively, if there are $\lambda$ virtual prompts that are $\epsilon$-close to the authentic one over $\mathcal{D}$ and $\epsilon$ is small, these virtual prompts appear authentic, as their probabilities are close to that of the authentic one, bounded by $\epsilon$. In the case that $\epsilon$ closes to zero, $P_{\mathcal{D}}(s)$ is about the same for all $s$ from these prompts. In other words, even if an adversary knows $\mathcal{D}$ and obtains the plaintext of all $\lambda+1$ prompts, it can only guess randomly, with a probability of approximately $\frac{1}{\lambda+1}$ of being correct, and an advantage bounded by $\epsilon$.
Worse Case Analysis  Theorem 5.2 implies that if both $\epsilon$ and $\Delta$ are small, the virtual prompts generated as described in §5.1 would appear as authentic as the user prompt $S$.
@@
Given a user prompt $S$ and a subsequence $S_{target}$, let $F=\text{GQS}(S,S_{\text{target}},\epsilon)$ and $\Delta=\Delta_{U}(\mathcal{D},\mathcal{D}_{\text{LM}})$, where $U=\{S_{target}\}\cup F$. Then $\{S_{target}\}$ and $F$ are $(\epsilon+2\Delta)$-close over $\mathcal{D}$.
More precisely, the maximum difference between the probabilities of any pair of authentic and fake subsequences is $\epsilon+2\Delta$, given $S$ as context. The proof is provided in the appendix. Suppose that we have an authentic prompt $S$ and a set of virtual prompts $S_{1},S_{2},\ldots,S_{\lambda}$. Let $\mathcal{D}_{\text{LM}}$ be the distribution of all possible sequences modeled by $LM$. Theorem 5.2 essentially states that an adversary with knowledge of $\mathcal{D}$ can guess the authentic prompt $S$ with a probability $\epsilon+2\Delta$ better than randomly choosing from $\lambda+1$ possible prompts. Therefore, the security of PO can be characterized by three factors: (1) the set size, $\lambda$; (2) the error bound of sampling, $\epsilon$; and (3) the gap between the true distribution and its approximation, $\Delta=\Delta_{U}(\mathcal{D},\mathcal{D}_{\text{LM}})$.
@@
However, PO does not fully guarantee confidentiality when the true distribution $\mathcal{D}$ differs from $\mathcal{D}_{\text{LM}}$. This can happen when $\mathcal{D}_{\text{LM}}$ is simply not sufficiently sophisticated. In such cases, PO serves as an obfuscation, as its name suggests, increases the cost of cryptanalysis, but does not fully prevent it. The above analysis suggests that an adversary with greater knowledge of the true distribution $\mathcal{D}$ gains an advantage in guessing. However, it is important to note that $\mathcal{D}$ is not always accessible to adversaries, meaning the actual error bound in practice is likely smaller than our theoretical analysis.
Another scenario in which PO does not provide full confidentiality is when the prompt $S$ is unique. In such cases, it may be impossible to generate fake n-grams without using a large $\epsilon$ or small $\lambda$, regardless of whether $\mathcal{D}$ is well captured by $\mathcal{D}_{\text{LM}}$ or not. To prevent a potential information leak, PO allows users to configure the desired $\epsilon$ and $\lambda_{min}$. If $\lambda<\lambda_{min}$, the CVM will stop in the prefill phase and alert the user.